find the solutions

2+2-x= ?


1
Expert's answer
2021-06-18T06:28:09-0400
"2+2-x=(2+2)-x=4-x"

If


"2+2-x=0"

Then


"4-x=0"

"4-x+x=0+x"

"4=x"

"x=4"